# Diagonal Simplicial Tensor Modules and Algebraic $n$-Hypergroupoids

### v1

@@ 8.4 Injectivity Analysis and Examples
The functor $I_{\bullet}:\Delta\to\mathbf{Set}$ preserves monomorphisms.
@@
Monomorphisms in $\Delta$ are generated by coface maps $\delta_{i}$. $I_{\bullet}(\delta_{i}^{p})=\Delta_{i}^{p}$ is a product of injections, hence injective. ∎
@@
    If $K^{\prime}_{0}>0$, the projection $\Psi_{0}:\mathcal{U}\to\operatorname{Gr}(K^{\prime}_{0},S_{0})$ is dominant.
@@
(2) Since $\mathcal{I}_{0}$ is injective, the entries of $M_{T_{\text{sym}},0}$ are distinct variables. The map $L_{0}:X_{n}\to\operatorname{Mat}_{R_{0}\times S_{0}}(K)$ sending $T$ to $M_{T,0}$ is a projection onto these coordinates, hence surjective. The map from the space of matrices of maximal rank $R^{\prime}_{0}$ to the Grassmannian of their kernels is dominant. $\Psi_{0}$ is the composition of $L_{0}$ (restricted to $\mathcal{U}$) and this dominant map. ∎
@@
$n=1,k=2$. $S=(2,3)$. $R=(1,4)$. $p=1$. $S_{1}=3,R_{1}=4$. The morphisms $\Delta([1],[1])$ are $\{\mathrm{id},(0,0),(1,1)\}$. $I_{1}=[1]^{2}$. The symbolic realization matrix $M_{T_{\text{sym}},1}$ is:
@@ 8.5 The Incidence Variety and Segre–Plücker Embedding
Embed $\operatorname{Gr}(\vec{s})$ into a projective space via the composition of the Plücker embeddings $\iota_{p}:\operatorname{Gr}(K^{\prime}_{p},S_{p})\hookrightarrow\mathbb{P}(\Lambda^{K^{\prime}_{p}}K[\Delta^{n}]_{p})$ and the Segre embedding $\Sigma$. The image $\Sigma(\mathbf{Gr}^{\mathrm{simp}}(\vec{s}))$ is the intersection of $\Sigma(\operatorname{Gr}(\vec{s}))$ with a linear subspace of the ambient projective space.
@@
We analyze the condition $d_{i}(K_{p})\subseteq K_{p-1}$. Let $C_{q}=K[\Delta^{n}]_{q}$. Let $\omega_{p}$ and $\omega_{p-1}$ be the Plücker coordinates representing $K_{p}$ and $K_{p-1}$. The condition is equivalent to $\Lambda^{K^{\prime}_{p}}d_{i}(\omega_{p})\wedge\eta_{p-1}=0$, where $\eta_{p-1}$ is a volume element of a complement to $K_{p-1}$. Under the Segre embedding, these bilinear equations in the Plücker coordinates become linear equations in the homogeneous coordinates of the tensor product space. ∎
@@
Let $C_{p}:=K[\Delta^{n}]_{p}$. On $\operatorname{Gr}(\vec{s})$, let $\mathcal{S}_{p}\subset C_{p}\otimes\mathcal{O}_{\operatorname{Gr}(\vec{s})}$ denote the pullback of the tautological rank-$K^{\prime}_{p}(\vec{s})$ subbundle on $\operatorname{Gr}(K^{\prime}_{p}(\vec{s}),C_{p})$ along the projection $\operatorname{Gr}(\vec{s})\to\operatorname{Gr}(K^{\prime}_{p}(\vec{s}),C_{p})$, and let $\mathcal{S}_{p-1}$ be defined similarly in degree p1p-1.
@@
A point of $\operatorname{Gr}(\vec{s})$ is a tuple of subspaces $(K_{p})_{p}$ with $K_{p}\subset C_{p}$ and $\dim K_{p}=K^{\prime}_{p}(\vec{s})$ whenever $K^{\prime}_{p}(\vec{s})>0$. By construction of the tautological subbundle, the fiber $(\mathcal{S}_{p})_{x}$ over a point $x$ corresponding to $(K_{p})_{p}$ is exactly $K_{p}$, and similarly $(\mathcal{S}_{p-1})_{x}=K_{p-1}$.
@@
$d_{i}\otimes\mathrm{id}:C_{p}\otimes\mathcal{O}_{\operatorname{Gr}(\vec{s})}\longrightarrow C_{p-1}\otimes\mathcal{O}_{\operatorname{Gr}(\vec{s})}.$
@@
At a point $x=(K_{p})_{p}$, the fiber map $(\varphi_{p,i})_{x}:K_{p}\to C_{p-1}/K_{p-1}$ is the composition
$K_{p}\hookrightarrow C_{p}\xrightarrow{d_{i}}C_{p-1}\twoheadrightarrow C_{p-1}/K_{p-1}.$
Thus $(\varphi_{p,i})_{x}=0$ if and only if $d_{i}(K_{p})\subseteq K_{p-1}$. By definition, $\mathbf{Gr}^{\mathrm{simp}}(\vec{s})$ is the locus of tuples satisfying these inclusions for all $(p,i)$.
@@ 8.6 Homology of Generated Subobjects
Fix $T\in X_{n}(\vec{s};K)$ and write $C_{\bullet}:=K[\Delta^{n}]$ and $K_{\bullet}:=K(T)_{\bullet}=\ker(f_{T})\subseteq C_{\bullet}$. We analyze $H_{\bullet}(\langle T\rangle_{\bullet})$ using the short exact sequence
@@ 8.7 Example: Shape $(3,3)$
    For $K[\Delta^{2}]_{1}$, the basis of monotone maps $1]\to[2]$ ordered as
@@
    For $K[\Delta^{2}]_{2}$, the basis of monotone maps $2]\to[2]$ ordered as the ten non-decreasing triples $(a_{0},a_{1},a_{2})$ with $0\leq a_{0}\leq a_{1}\leq a_{2}\leq 2$:
@@
Two obvious kernel vectors can be written purely in terms of the diagonal entries. Set
@@
$M_{T,2}=\begin{pmatrix}v_{00}&v_{00}&v_{00}&v_{11}&v_{11}&v_{22}&v_{11}&v_{11}&v_{22}&v_{22}\\ v_{00}&v_{00}&v_{00}&v_{01}&v_{01}&v_{02}&v_{11}&v_{11}&v_{12}&v_{22}\\ v_{00}&v_{00}&v_{00}&v_{10}&v_{10}&v_{20}&v_{11}&v_{11}&v_{21}&v_{22}\\ v_{00}&v_{00}&v_{01}&v_{11}&v_{12}&v_{22}&v_{11}&v_{12}&v_{22}&v_{22}\\ v_{00}&v_{01}&v_{02}&v_{11}&v_{12}&v_{22}&v_{11}&v_{12}&v_{22}&v_{22}\\ v_{00}&v_{02}&v_{02}&v_{12}&v_{12}&v_{22}&v_{12}&v_{12}&v_{22}&v_{22}\\ v_{00}&v_{00}&v_{10}&v_{11}&v_{21}&v_{22}&v_{11}&v_{21}&v_{22}&v_{22}\\ v_{00}&v_{10}&v_{20}&v_{11}&v_{21}&v_{22}&v_{11}&v_{21}&v_{22}&v_{22}\\ v_{00}&v_{20}&v_{20}&v_{21}&v_{21}&v_{22}&v_{21}&v_{21}&v_{22}&v_{22}\end{pmatrix}$.
@@ A.1 The Chain Complex and Homotopy Operator
We verify the chain contraction identity $\partial_{p+1}H_{p}+H_{p-1}\partial_{p}=\mathrm{id}$. Using Lemma A.3 (valid for all $p\geq 0$):
$\partial_{p+1}H_{p}$ $=d_{0}H_{p}+\sum_{i=1}^{p+1}(-1)^{i}d_{i}H_{p}$
$=\mathrm{id}+\sum_{i=1}^{p+1}(-1)^{i}H_{p-1}d_{i-1}^{p}.$
Re-indexing the sum (j=i1j=i-1):
$\sum_{i=1}^{p+1}(-1)^{i}H_{p-1}d_{i-1}^{p}$ $=\sum_{j=0}^{p}(-1)^{j+1}H_{p-1}d_{j}^{p}=-\,H_{p-1}\partial_{p}.$
(This holds even for $p=0$, as $\partial_{0}=0$ and $H_{-1}=0$). Substituting back, we obtain
$\partial_{p+1}H_{p}+H_{p-1}\partial_{p}\;=\;\mathrm{id}_{X_{p}},$
so $H$ is a chain contraction of $X_{\bullet}(\vec{s};A)$. ∎
@@ A.2 Equivariance Properties of the Homotopy
The definition of $H_{p}$ depends on the presence of zero coordinates. Since permutation only changes the positions of the coordinates, the condition $\exists a:m_{a}=0$” is equivalent to $\exists b:(\sigma^{-1}m)_{b}=0$”.
@@
$H_{p}(\sigma\cdot T)(m)$ $=(\sigma\cdot T)(m-\vec{1})$
$=T(\sigma^{-1}(m-\vec{1})).$
@@
$(\sigma\cdot H_{p}(T))(m)$ $=H_{p}(T)(\sigma^{-1}m)$
$=T((\sigma^{-1}m)-\vec{1}).$
@@ A.3 Shifted Depth Filtration and Spectral Sequence
The differential $\partial$ and the homotopy $H$ preserve the shifted depth filtration (filtration degree 0). For all $p\geq 0$ and $t\in\mathbb{Z}$:
@@
    $d_{i}^{p}(\mathcal{F}^{t}X_{p})\subseteq\mathcal{F}^{t}X_{p-1}$ for all $i\in\{0,\dots,p\}$.
@@
(1) Let $T\in\mathcal{F}^{t}X_{p}$. We want to show $d_{i}^{p}(T)\in\mathcal{F}^{t}X_{p-1}$. Let $m\in I_{p-1}$. We must show that if $\min(m)<t+p-1$, then $d_{i}^{p}(T)(m)=0$. $d_{i}^{p}(T)(m)=T(\Delta_{i}^{p}(m))$. The coface maps satisfy $\delta_{i}(x)\leq x+1$ for all $i,x$. Thus $\min(\Delta_{i}^{p}(m))\leq\min(m)+1$. If $\min(m)<t+p-1$, then $\min(\Delta_{i}^{p}(m))<t+p$. Since $T\in\mathcal{F}^{t}X_{p}$, $T(\Delta_{i}^{p}(m))=0$. Thus $d_{i}^{p}(T)\in\mathcal{F}^{t}X_{p-1}$.
(2) Let $T\in\mathcal{F}^{t}X_{p}$. We want to show $H_{p}(T)\in\mathcal{F}^{t}X_{p+1}$. Let $m\in I_{p+1}$. We must show that if $\min(m)<t+p+1$, then $H_{p}(T)(m)=0$.
If $\min(m)=0$, $H_{p}(T)(m)=0$ by definition. If $\min(m)>0$, $H_{p}(T)(m)=T(m-\vec{1})$. If $\min(m)<t+p+1$, then $\min(m-\vec{1})=\min(m)-1<t+p$. Since $T\in\mathcal{F}^{t}X_{p}$, $T(m-\vec{1})=0$. Thus $H_{p}(T)\in\mathcal{F}^{t}X_{p+1}$. ∎
@@
Hence it collapses at the $E_{1}$–-page.
@@
By Lemma A.7, $(X_{\bullet},\partial,\mathcal{F}^{\bullet})$ is a filtered complex, as $\partial$ has filtration degree 0. Let $E_{0}=\operatorname{gr}_{\mathcal{F}}(X_{\bullet})$ be the associated graded complex, with differential $d^{0}$ induced by $\partial$.
By Lemma A.7, the homotopy operator $H$ also has filtration degree 0. It induces a map $[H]$ on $E_{0}$. The chain contraction identity $\partial H+H\partial=\mathrm{id}$ holds in the filtered complex. Since all operators involved preserve the filtration, the identity descends to the associated graded complex:
$d^{0}[H]+[H]d^{0}=[\mathrm{id}]=\mathrm{id}_{E_{0}}.$

### v2

@@ 8.4 Injectivity Analysis and Examples
The functor $I_{\bullet}:\boldsymbol{\Delta}\to\mathbf{Set}$ preserves monomorphisms.
@@
Monomorphisms in $\boldsymbol{\Delta}$ are generated by coface maps $\delta_{i}$. $I_{\bullet}(\delta_{i}^{p})=\Delta_{i}^{p}$ is a product of injections, hence injective. ∎
@@
    If $K^{\prime}_{0}>0$, the projection $\Psi_{0}:\mathcal{U}\to\operatorname{Gr}(K^{\prime}_{0},K[\boldsymbol{\Delta}^{n}]_{0})$ is dominant.
@@
(2) Since $\mathcal{I}_{0}$ is injective, the entries of $M_{T_{\mathrm{sym}},0}$ are distinct variables. The map $L_{0}:X_{n}\to\operatorname{Mat}_{R_{0}\times S_{0}}(K)$ sending $T$ to $M_{T,0}$ is a projection onto these coordinates, hence surjective. The map from the space of matrices of maximal rank $R^{\prime}_{0}$ to the Grassmannian of their kernels is dominant. $\Psi_{0}$ is the composition of $L_{0}$ (restricted to $\mathcal{U}$) and this dominant map. ∎
@@
$n=1,k=2$. $S=(2,3)$. $R=(1,4)$. $p=1$. $S_{1}=3,R_{1}=4$. The morphisms $\boldsymbol{\Delta}([1],[1])$ are $\{\operatorname{id},(0,0),(1,1)\}$. $I_{1}=[1]^{2}$. The symbolic realization matrix $M_{T_{\mathrm{sym}},1}$ is:
@@ 8.5 The Incidence Variety and Segre–Plücker Embedding
Embed $\operatorname{Gr}(\vec{s})$ into a projective space via the composition of the Plücker embeddings $\iota_{p}:\operatorname{Gr}(K^{\prime}_{p},S_{p})\hookrightarrow\mathbb{P}(\Lambda^{K^{\prime}_{p}}K[\boldsymbol{\Delta}^{n}]_{p})$ and the Segre embedding $\Sigma$. The image $\Sigma(\mathbf{Gr}^{\mathrm{simp}}(\vec{s}))$ is the intersection of $\Sigma(\operatorname{Gr}(\vec{s}))$ with a linear subspace of the ambient projective space.
@@
We analyze the condition $d_{i}(L_{p})\subseteq L_{p-1}$. Let $C_{q}=K[\boldsymbol{\Delta}^{n}]_{q}$. Let $\omega_{p}$ and $\omega_{p-1}$ be the Plücker coordinates representing $L_{p}$ and $L_{p-1}$. The condition is equivalent to $\Lambda^{K^{\prime}_{p}}d_{i}(\omega_{p})\wedge\eta_{p-1}=0$, where $\eta_{p-1}$ is a volume element of a complement to $L_{p-1}$. Under the Segre embedding, these bilinear equations in the Plücker coordinates become linear equations in the homogeneous coordinates of the tensor product space. ∎
@@
Let $C_{p}:=K[\boldsymbol{\Delta}^{n}]_{p}$. On $\operatorname{Gr}(\vec{s})$, let $\mathcal{S}_{p}\subset C_{p}\otimes\mathcal{O}_{\operatorname{Gr}(\vec{s})}$ denote the pullback of the tautological rank-$K^{\prime}_{p}(\vec{s})$ subbundle on $\operatorname{Gr}(K^{\prime}_{p}(\vec{s}),C_{p})$ along the projection $\operatorname{Gr}(\vec{s})\to\operatorname{Gr}(K^{\prime}_{p}(\vec{s}),C_{p})$, and let $\mathcal{S}_{p-1}$ be defined similarly in degree p1p-1.
@@
A point of $\operatorname{Gr}(\vec{s})$ is a tuple of subspaces $(L_{p})_{p}$ with $L_{p}\subset C_{p}$ and $\dim L_{p}=K^{\prime}_{p}(\vec{s})$ whenever $K^{\prime}_{p}(\vec{s})>0$. By construction of the tautological subbundle, the fiber $(\mathcal{S}_{p})_{x}$ over a point $x$ corresponding to $(L_{p})_{p}$ is exactly $L_{p}$, and similarly $(\mathcal{S}_{p-1})_{x}=L_{p-1}$.
@@
$d_{i}\otimes\operatorname{id}:C_{p}\otimes\mathcal{O}_{\operatorname{Gr}(\vec{s})}\longrightarrow C_{p-1}\otimes\mathcal{O}_{\operatorname{Gr}(\vec{s})}.$
@@
At a point $x=(L_{p})_{p}$, the fiber map $(\varphi_{p,i})_{x}:L_{p}\to C_{p-1}/L_{p-1}$ is the composition
$L_{p}\hookrightarrow C_{p}\xrightarrow{d_{i}}C_{p-1}\twoheadrightarrow C_{p-1}/L_{p-1}.$
Thus $(\varphi_{p,i})_{x}=0$ if and only if $d_{i}(L_{p})\subseteq L_{p-1}$. By definition, $\mathbf{Gr}^{\mathrm{simp}}(\vec{s})$ is the locus of tuples satisfying these inclusions for all $(p,i)$.
@@ 8.6 Homology of Generated Subobjects
Fix $T\in X_{n}(\vec{s};K)$ and write $C_{\bullet}:=K[\boldsymbol{\Delta}^{n}]$ and $K_{\bullet}:=K(T)_{\bullet}=\ker(f_{T})\subseteq C_{\bullet}$. We analyze $H_{\bullet}(\langle T\rangle_{\bullet})$ using the short exact sequence
@@ 8.7 Example: Shape $(3,3)$
    For $K[\boldsymbol{\Delta}^{2}]_{1}$, the basis of monotone maps $1]\to[2]$ ordered as
@@
    For $K[\boldsymbol{\Delta}^{2}]_{2}$, the basis of monotone maps $2]\to[2]$ ordered as the ten non-decreasing triples $(a_{0},a_{1},a_{2})$ with $0\leq a_{0}\leq a_{1}\leq a_{2}\leq 2$:
@@
Two kernel vectors can be written in terms of the diagonal entries. Set
@@
$M_{T,2}=\begin{pmatrix}v_{00}&v_{00}&v_{00}&v_{00}&v_{00}&v_{00}&v_{11}&v_{11}&v_{11}&v_{22}\\ v_{00}&v_{00}&v_{00}&v_{01}&v_{01}&v_{02}&v_{11}&v_{11}&v_{12}&v_{22}\\ v_{00}&v_{01}&v_{02}&v_{01}&v_{02}&v_{02}&v_{11}&v_{12}&v_{12}&v_{22}\\ v_{00}&v_{00}&v_{00}&v_{10}&v_{10}&v_{20}&v_{11}&v_{11}&v_{21}&v_{22}\\ v_{00}&v_{00}&v_{00}&v_{11}&v_{11}&v_{22}&v_{11}&v_{11}&v_{22}&v_{22}\\ v_{00}&v_{01}&v_{02}&v_{11}&v_{12}&v_{22}&v_{11}&v_{12}&v_{22}&v_{22}\\ v_{00}&v_{10}&v_{20}&v_{10}&v_{20}&v_{20}&v_{11}&v_{21}&v_{21}&v_{22}\\ v_{00}&v_{10}&v_{20}&v_{11}&v_{21}&v_{22}&v_{11}&v_{21}&v_{22}&v_{22}\\ v_{00}&v_{11}&v_{22}&v_{11}&v_{22}&v_{22}&v_{11}&v_{22}&v_{22}&v_{22}\end{pmatrix}$.
@@ A.1 The Chain Complex and Homotopy Operator
We verify the chain contraction identity
$\partial_{p+1}H_{p}+H_{p-1}\partial_{p}=\operatorname{id}_{X_{p}}\qquad(p\geq 0).$
Base case $p=0$. Here $\partial_{0}=0$ and $H_{-1}=0$. Also $\partial_{1}=d_{0}^{1}-d_{1}^{1}$. Thus
$\partial_{1}H_{0}=(d_{0}^{1}-d_{1}^{1})H_{0}=d_{0}^{1}H_{0}-d_{1}^{1}H_{0}=\operatorname{id}_{X_{0}}-0=\operatorname{id}_{X_{0}},$
using Lemma A.3(1) for $p=0$ and Lemma A.3(2) in the special $p=0$ verification (i.e. $d_{1}^{1}H_{0}=0$). Therefore
$\partial_{1}H_{0}+H_{-1}\partial_{0}=\operatorname{id}_{X_{0}}.$
Inductive case $p\geq 1$. Using Lemma A.3:
$\partial_{p+1}H_{p}$ $=d_{0}^{p+1}H_{p}+\sum_{i=1}^{p+1}(-1)^{i}d_{i}^{p+1}H_{p}$
$=\operatorname{id}_{X_{p}}+\sum_{i=1}^{p+1}(-1)^{i}H_{p-1}d_{i-1}^{p}.$
Re-indexing with j=i1j=i-1 gives
$\sum_{i=1}^{p+1}(-1)^{i}H_{p-1}d_{i-1}^{p}$ $=\sum_{j=0}^{p}(-1)^{j+1}H_{p-1}d_{j}^{p}$
$=-\,H_{p-1}\partial_{p}.$
Substituting back yields
$\partial_{p+1}H_{p}+H_{p-1}\partial_{p}=\operatorname{id}_{X_{p}}.$
Thus $H$ is a chain contraction of $X_{\bullet}(\vec{s};A)$. ∎
@@ A.2 Equivariance Properties of the Homotopy
The definition of $H_{p}$ depends on the presence of zero coordinates.
@@
$H_{p}(\sigma\cdot T)(m)$ $=(\sigma\cdot T)(m-\vec{1})=T(\sigma^{-1}(m-\vec{1})).$
@@
$(\sigma\cdot H_{p}(T))(m)$ $=H_{p}(T)(\sigma^{-1}m)=T((\sigma^{-1}m)-\vec{1}).$
@@ A.3 Shifted Depth Filtration and Spectral Sequence
The differential $\partial$ and the homotopy $H$ preserve the shifted depth filtration (filtration degree 0). For all $t\in\mathbb{Z}$:
@@
    For all $p\geq 1$ and all $i\in\{0,\dots,p\}$,
    $d_{i}^{p}(\mathcal{F}^{t}X_{p})\subseteq\mathcal{F}^{t}X_{p-1}.$
@@
(1) Let $p\geq 1$ and let $T\in\mathcal{F}^{t}X_{p}$. We show $d_{i}^{p}(T)\in\mathcal{F}^{t}X_{p-1}$. Let $m\in I_{p-1}$. If $\min(m)<t+p-1$, then
$d_{i}^{p}(T)(m)=T(\Delta_{i}^{p}(m)).$
For each coordinate, the coface map satisfies $\delta_{i}(x)\leq x+1$, hence
$\min(\Delta_{i}^{p}(m))\leq\min(m)+1.$
If $\min(m)<t+p-1$, then $\min(\Delta_{i}^{p}(m))<t+p$, and since $T\in\mathcal{F}^{t}X_{p}$ we have $T(\Delta_{i}^{p}(m))=0$. Therefore $d_{i}^{p}(T)\in\mathcal{F}^{t}X_{p-1}$.
(2) Let $p\geq 0$ and let $T\in\mathcal{F}^{t}X_{p}$. We show $H_{p}(T)\in\mathcal{F}^{t}X_{p+1}$. Let $m\in I_{p+1}$. If $\min(m)=0$, then $H_{p}(T)(m)=0$ by definition of $H$. If $\min(m)>0$, then $H_{p}(T)(m)=T(m-\vec{1})$. If $\min(m)<t+p+1$, then
$\min(m-\vec{1})=\min(m)-1<t+p,$
so $T(m-\vec{1})=0$ because $T\in\mathcal{F}^{t}X_{p}$. Hence $H_{p}(T)\in\mathcal{F}^{t}X_{p+1}$. ∎
@@
Hence it collapses at the $E_{1}$-page.
@@
By Lemma A.7, for $p\geq 1$ the differential $\partial_{p}$ has filtration degree 0, and since $\partial_{0}=0$ the filtered complex $(X_{\bullet},\partial,\mathcal{F}^{\bullet})$ is well-defined.
Let $E_{0}=\operatorname{gr}_{\mathcal{F}}(X_{\bullet})$ be the associated graded complex, with differential $d^{0}$ induced by $\partial$.
By Lemma A.7, the homotopy operator $H$ also has filtration degree 0. It induces a map $[H]$ on $E_{0}$. The chain contraction identity $\partial H+H\partial=\operatorname{id}$ holds in the filtered complex. Since all operators involved preserve the filtration, the identity descends to the associated graded complex:
$d^{0}[H]+[H]d^{0}=[\operatorname{id}]=\operatorname{id}_{E_{0}}.$